Calculate Å B ☺ Ay 0 B By
Answer #1

It seems the problem és like this TA J3 90 Y00 a - A+ AS = AG3 + Afn3003 A의 = 층서 불3 82 ㅎㅎ B처 83 -860% + Bugs =B BE B = I N E